# Kernel Ridge Riesz Representers: Generalization Error and Mis-specification

### v3

@@ 3.3 RKHS notation and key assumptions
Similar to kernel ridge regression, the kernel ridge balancing weight algorithm conducts estimation in a reproducing kernel Hilbert space (RKHS) $H\subset L_{2}$. $H$ consists of functions of the form $f:\mathcal{W}\rightarrow\mathbb{R}$. I denote its symmetric, positive definite kernel by $k:\mathcal{W}\times\mathcal{W}\rightarrow\mathbb{R}$, and its feature map by $\phi:w\mapsto k(w,\cdot)$, so that $k(w,w^{\prime})=\langle\phi(w),\phi(w^{\prime})\rangle_{H}$ and $f(w)=\langle f,\phi(w)\rangle_{H}$ for $f\in H$ [^3] Assumptions 4.1 and 4.2 below formalize weak regularity conditions.
I use familiar spectral notation from kernel ridge regression analysis. Let $L:L_{2}\rightarrow L_{2}$ be the convolution operator $f\mapsto\int k(\cdot,w)f(w)\mathrm{d}\mathbb{P}(w)$. Since $L$ is self adjoint and positive definite, it has weakly decreasing countable eigenvalues $(\eta_{j})$ and corresponding eigenfunctions $(\varphi_{j})$. Define the space $H^{c}=(f=\sum_{j=1}^{\infty}f_{j}\varphi_{j}:\sum_{j=1}^{\infty}f_{j}^{2}\eta_{j}^{-c}<\infty)$. In this notation, $H^{0}=L_{2}$ and $H^{1}=H$ when $k$ is characteristic [Sriperumbudur et al., 2010]. In words, the RKHS $H$ is the subset of $L_{2}$ for which higher order terms in the series $(\varphi_{j})$ have a smaller contribution. By Mercer’s theorem, the feature map can be expressed as $\phi(w)=\{\eta_{j}^{1/2}\varphi_{j}(w)\}^{\infty}_{j=1}$.
@@
I place two standard assumptions from kernel ridge regression analysis. The first, called the source condition, concerns an object I call $\alpha_{H}\in H$, which is the best RKHS approximation to the Riesz representer $\alpha_{0}\in L_{2}$. [^4] For ATE, $\alpha_{H}\in\operatorname*{\arg\!\min}_{\alpha\in H}\|\alpha-\alpha_{0}\|_{2}$ and $\alpha_{0}(W)=D\pi_{0}(X)^{-1}-(1-D)\{1-\pi_{0}(X)\}^{-1}$. This assumption helps to analyze bias. The second assumption is the spectral decay condition, which quantifies the effective dimension of the features $\phi(W)$ of the RKHS $H$, and helps to analyze variance. Following [Caponnetto and De Vito, 2007], I denote by $\mathcal{P}(b,c)$ the class of distributions that satisfy these two assumptions.
Assumption 3.2 (Smoothness).
@@
By definition of $\alpha_{H}\in H$ as the best kernel approximation to $\alpha_{0}\in L_{2}$, $c\geq 1$ automatically holds in Assumption 3.2. A value $c>1$ means that $\alpha_{H}$ is a particularly smooth element of $H$. The $L_{2}$ rates in this paper do not improve beyond $c=2$, reflecting the well known saturation effect of Tikhonov regularization [Bauer et al., 2007].
Assumption 3.3 (Spectral decay).
@@
For any bounded kernel $k$, $b\geq 1$ automatically holds in Assumption 3.3(i) [Fischer and Steinwart, 2020, Lemma 10]. A value $b>1$ means that the RKHS has a lower effective dimension in light of the data distribution. The $L_{2}$ rates in this paper improve all the way to the parameteric rate as $b\rightarrow\infty$.
The upper bound, Assumption 3.3(i), applies to the Matérn, Gaussian, and other kernels with spectral decay that is polynomial, exponential, or “better”. The lower bound, Assumption 3.3(ii), applies to the Matérn kernel but not the Gaussian kernel. The main results use Assumption 3.3(i) only; Assumption 3.3(ii) is for comparison to known optimality theory.
@@ 3.4 Preview of main results
Lemma 3.2 (Main lemma).
Under Assumption 3.1 and weak regularity conditions (Assumptions 4.1 and 4.2 below), the counterfactual effective dimension of $\phi^{(m)}(W)$ is upper bounded by $\bar{M}$ times the effective dimension of $\phi(W)$. See Section 6 for details.
By Lemma 3.2, a familiar regularity condition from semiparametric theory (Assumption 3.1), together with a familiar condition from learning theory that controls the effective dimension (Assumption 3.3), implies control of the counterfactual effective dimension and hence the variance of $\tilde{\alpha}$. Due to this insight, I avoid an additional approximation assumption on the counterfactual effective dimension when proving the main result; see Remark 6.1.
Theorem 3.1 (Main theoretical result).
Under Assumptions 3.13.2 and 3.3(i) as well as weak regularity conditions (Assumptions 4.1 and 4.2 below), $\|\tilde{\alpha}-\alpha_{H}\|_{2}^{2}=O_{p}\{n^{-bc/(bc+1)}\}$ using regularization $\lambda=n^{-b/(bc+1)}$, when $b\in(1,\infty)$, $c\in(1,2]$, and $\bar{M}$ is fixed. This rate is optimal for some choices of $m$ when Assumption 3.3(ii) also holds. See Section 5 for details, including results when $b=\infty$, $c=1$, or $\bar{M}_{h}\uparrow\infty$.
Corollary 3.1 (Main corollary).
Under weak regularity conditions (stated below) and correct specification of $\hat{\gamma}$ and $\hat{\alpha}$, perhaps outside of the RKHS, if $\bar{M}$ is fixed then $\hat{\theta}$ converges to $\theta_{0}$ at the rate $n^{-1/2}$ and is asymptotically normal. Under mis-specification of $\hat{\gamma}$ or $\hat{\alpha}$, $\hat{\theta}$ converges to $\theta_{0}$, albeit at a slower rate. For pointwise approximations of nonparametric causal functions with $\bar{M}_{h}\uparrow\infty$, $\hat{\theta}$ converges to $\theta_{0}$ at the rate $(nh)^{-1/2}$ and is asymptotically normal when the bandwidth satisfies $h=o(1)$ and $n^{-1/2}h^{-3/2}=o(1)$. See Section 5 and Appendix A for details.
Proposition 3.1 (A practical result).
@@ 4.1 General loss function
An important departure from many works on kernel ridge balancing weights is that I allow incorrect specification by the RKHS. [^5] See Remark 5.3 below. In particular, I allow $\gamma_{0}\not\in H$ and $\alpha_{0}\not\in H$. For the scenario where $\alpha_{0}\not\in H$, I define the best kernel approximation $\alpha_{H}\in H$ to the Riesz representer $\alpha_{0}\in L_{2}$. I defer proofs for this section to Appendix B.
Definition 4.1 (Kernel approximation to Riesz representer).
@@
Assumption 4.1 (Original spaces).
@@
Assumption 4.2 (RKHS regularity).
@@
The space $\mathcal{W}$ accommodates general data types, e.g. texts, images, and graphs. Commonly used kernels are bounded. Boundedness implies that the counterfactual kernel mean embedding $\mu^{(m)}=\mathbb{E}\{\phi^{(m)}(W)\}$ and the uncentered covariance operator $T=\mathbb{E}\{\phi(W)\otimes\phi(W)^{*}\}$ are well defined. [^6] Here, I use outer product notation: $(a\otimes b^{*})c=a\langle b,c\rangle_{H}$, where $b^{*}$ is the adjoint of $b$. The counterfactual kernel mean embedding generalizes the well known kernel mean embedding [Smola et al., 2007]. It is unconditional and does not involve the outcome, unlike outcome embeddings studied by e.g. [Muandet et al., 2021, Singh et al., 2024]. Measurability is another standard condition.
The characteristic property ensures that no counterfactual information is lost when approximating $\alpha_{0}$ as $\alpha_{H}$. In particular, the characteristic property means that $\mathbb{P}\mapsto\mu^{(m)}$ is injective, so $\mu^{(m)}$ may be used as a sufficient statistic for $\alpha_{H}$, as shown in the following lemma.
Lemma 4.1 (Towards a loss for KRRR).
Suppose Assumptions 3.14.1, and 4.2 hold. Then we have $\alpha_{H}\in\operatorname*{\arg\!\min}_{\alpha\in H}\mathcal{L}(\alpha)$ where $\mathcal{L}(\alpha)=-2\langle\alpha,\mu^{(m)}\rangle_{H}+\langle\alpha,T\alpha\rangle_{H}$.
The sufficient statistics $\mu^{(m)}\in H$ and $T\in H\otimes H^{*}$ are infinite dimensional generalizations of the counterfactual moment and the covariance matrix in automatic debiased machine learning (Auto-DML) [Chernozhukov et al., 2022d, Chernozhukov et al., 2022c]. Auto-DML uses $p$ explicit basis functions, and its sufficient statistic are a vector in $\mathbb{R}^{p}$ and a matrix in $\mathbb{R}^{p\times p}$. Kernel balancing weights use the countable eigenfunctions of the kernel $k$ as infinitely many implicit basis functions.
Inspired by Lemma 4.1, I study a general loss for kernel ridge Riesz representers (KRRR) at the population level. As we will see below, its empirical analogue recovers the loss for kernel ridge balancing weights previously proposed for treatment effects. Let $\hat{\mu}^{(m)}=\mathbb{E}_{n}\{\phi^{(m)}(W)\}$ and $\hat{T}=\mathbb{E}_{n}\{\phi(W)\otimes\phi(W)^{*}\}$ be the empirical summary statistics.
Definition 4.2 (Loss for KRRR).
@@
The KRRR loss in Definition 4.2 differs from Dantzig [Chernozhukov et al., 2022d], Lasso [Chernozhukov et al., 2022c], and adversarial [Chernozhukov et al., 2020] losses for Riesz representers whose generalization errors have been previously analyzed. The KRRR loss has a simple first order condition which implies a balance-variance trade-off similar to those of kernel balancing weights in the literature [Hazlett, 2020, Wong and Chan, 2018, Zhao, 2019, Kallus, 2020, Hirshberg et al., 2019].
Lemma 4.2 (First order condition).
Suppose the conditions of Lemma 4.1 hold. The first order conditions give $T\alpha_{H}=\mu^{(m)}$, $\alpha_{\lambda}=(T+\lambda)^{-1}\mu^{(m)}$, and $\tilde{\alpha}=(\hat{T}+\lambda)^{-1}\hat{\mu}^{(m)}.$
Corollary 4.1 (Balance-variance trade-off).
Suppose the conditions of Lemma 4.1 hold. Then we have $\|\mathbb{E}_{n}\{\tilde{\alpha}(W)\phi(W)-\phi^{(m)}(W)\}\|_{H}/\|\tilde{\alpha}\|_{H}=\lambda$.
By Definition 4.2, a smaller regularization value $\lambda$ translates to an estimator $\tilde{\alpha}$ with a smaller bias and a larger variance. Corollary 4.1 illustrates that a smaller (normalized) imbalance of the features, $\|\mathbb{E}_{n}[\tilde{\alpha}(W)\phi(W)-\phi^{(m)}(W)]\|_{H}/\|\tilde{\alpha}\|_{H}$, is possible if and only if we pay the cost of a larger variance. In the finite sample analysis to follow, $\lambda\downarrow 0$ in a way that optimally navigates this trade-off (when rate lower bounds are known). In particular, we will not require undersmoothing for inference on the causal parameter.
4.2 Detailed comparisons to kernel ridge regression and kernel ridge balancing weights
Lemma 4.2 has several more corollaries, which formalize the sense in which KRRR builds on and extends previous frameworks. In these corollaries, let $\tilde{\gamma}$ be the kernel ridge regression estimator. Throughout, I assume that the conditions of Lemma 4.1 hold.
Corollary 4.2 (Special case: Kernel ridge regression).
@@ 4.2 Detailed comparisons to kernel ridge regression and kernel ridge balancing weights
Definition 4.3 (Loss for kernel ridge balancing weights).
@@
Corollary 4.3 (Special case: Kernel ridge balancing weights; eq. 8 of [Bruns-Smith et al., 2023]).
On the training data, $\tilde{\alpha}(W_{i})=\tilde{\beta}_{i}$. Appendix D.3 of [Hirshberg et al., 2019] studies $\frac{1}{n}\sum_{i=1}^{n}\{\tilde{\beta}_{i}-\alpha_{0}(D_{i},X_{i})\}^{2}=\mathbb{E}_{n}[\{\tilde{\alpha}(D,X)-\alpha_{0}(D,X)\}^{2}]$, i.e. training error.
In summary, KRRR nests kernel ridge regression and kernel ridge balancing weights, unlike the adversarial Riesz representer of [Chernozhukov et al., 2020]. The connection to kernel ridge regression foreshadows the generalization error guarantees for KRRR which are the focus of this work.
Next, I recap known equivalences for the estimation of treatment effects [Speckman, 1979, Kallus, 2020, Hirshberg et al., 2019] and more generally for the estimation of causal parameters within the class of Assumption 3.1 [Bruns-Smith and Feller, 2022, Bruns-Smith et al., 2023]. With $\hat{\alpha}$ taken to be KRRR $\tilde{\alpha}$, these equivalences only hold when the regression estimator $\hat{\gamma}$ is kernel ridge regression $\tilde{\gamma}$, and when the same observations are used for $\tilde{\gamma}$, $\tilde{\alpha}$, and the causal parameter.
Corollary 4.4 (Equivalence in a special case; Proposition 27 of [Kallus, 2020], Lemma 1 of [Hirshberg et al., 2019], eq. 23 of [Bruns-Smith et al., 2023]).
@@
Corollary 4.5 (Non-equivalence in general).

### v4

@@ 3.3   RKHS notation and key assumptions
The kernel ridge balancing weight algorithm conducts estimation in a reproducing kernel Hilbert space (RKHS) $H\subset L_{2}$. $H$ consists of functions of the form $f:\mathcal{W}\rightarrow\mathbb{R}$. I denote its symmetric, positive definite kernel by $k:\mathcal{W}\times\mathcal{W}\rightarrow\mathbb{R}$, and its feature map by $\phi:w\mapsto k(w,\cdot)$, so that $k(w,w^{\prime})=\langle\phi(w),\phi(w^{\prime})\rangle_{H}$ and $f(w)=\langle f,\phi(w)\rangle_{H}$ for $f\in H$ [^4] Assumptions 4 and 5 below formalize weak regularity conditions.
I use familiar spectral notation from kernel ridge regression analysis. Let $L:L_{2}\rightarrow L_{2}$ be the convolution operator $f\mapsto\int k(\cdot,w)f(w)\mathrm{d}\mathbb{P}(w)$. Since $L$ is self adjoint and positive definite, it has weakly decreasing countable eigenvalues $(\eta_{j})$ and corresponding eigenfunctions $(\varphi_{j})$. Define the space $H^{c}=(f=\sum_{j=1}^{\infty}f_{j}\varphi_{j}:\sum_{j=1}^{\infty}f_{j}^{2}\eta_{j}^{-c}<\infty)$. In this notation, $H^{0}=L_{2}$ and $H^{1}=H$ when $k$ is characteristic [Sriperumbudur et al., 2010]. The RKHS $H$ is the subset of $L_{2}$ for which higher order terms in the series $(\varphi_{j})$ have a smaller contribution. By Mercer’s theorem, the feature map is $\phi(w)=\{\eta_{j}^{1/2}\varphi_{j}(w)\}^{\infty}_{j=1}$.
@@
I place two standard assumptions from kernel ridge regression analysis. The first, called the source condition, concerns an object I call $\alpha_{H}\in H$, which is the best RKHS approximation to the Riesz representer $\alpha_{0}\in L_{2}$. [^5] For ATE, $\alpha_{H}\in\operatorname*{\arg\!\min}_{\alpha\in H}\|\alpha-\alpha_{0}\|_{2}$ and $\alpha_{0}(W)=D\pi_{0}(X)^{-1}-(1-D)\{1-\pi_{0}(X)\}^{-1}$. This assumption helps to analyze bias. The second assumption is the spectral decay condition, which quantifies the effective dimension of the features $\phi(W)$ of the RKHS $H$, and helps to analyze variance. I denote by $\mathcal{P}(b,c)$ the class of distributions that satisfy these two assumptions.
Assumption 2 (Smoothness).
@@
By definition of $\alpha_{H}\in H$ as the best kernel approximation to $\alpha_{0}\in L_{2}$, $c\geq 1$ automatically holds in Assumption 2. A value $c>1$ means that $\alpha_{H}$ is a particularly smooth element of $H$. The $L_{2}$ rates in this paper do not improve beyond $c=2$, reflecting the well known saturation effect of Tikhonov regularization [Bauer et al., 2007].
Assumption 3 (Spectral decay).
@@
For any bounded kernel $k$, $b\geq 1$ automatically holds in Assumption 3(i) [Fischer and Steinwart, 2020, Lemma 10]. A value $b>1$ means that the RKHS has a lower effective dimension in light of the data distribution. The $L_{2}$ rates in this paper improve all the way to the parameteric rate as $b\rightarrow\infty$.
The upper bound, Assumption 3(i), applies to the Matérn, Gaussian, and other kernels with spectral decay that is polynomial, exponential, or “better”. The lower bound, Assumption 3(ii), applies to the Matérn kernel but not the Gaussian kernel. The main results use Assumption 3(i) only; Assumption 3(ii) is for comparison to optimality theory.
@@ 3.4   Preview of main results
Lemma 2 (Main lemma).
Under Assumption 1 and weak regularity conditions (Assumptions 4 and 5 below), the counterfactual effective dimension of $\phi^{(m)}(W)$ is upper bounded by $\bar{M}$ times the effective dimension of $\phi(W)$. See Section 6 for details.
By Lemma 2, a familiar regularity condition from semiparametric theory (Assumption 1), together with a familiar effective dimension condition from learning theory (Assumption 3), implies control of the counterfactual effective dimension and hence the variance of $\tilde{\alpha}$. Due to this insight, I avoid an additional approximation assumption on the counterfactual effective dimension when proving the main result; see Remark 5.
Theorem 1 (Main theoretical result).
Under Assumptions 12 and 3(i) as well as weak regularity conditions (Assumptions 4 and 5 below), $\|\tilde{\alpha}-\alpha_{H}\|_{2}^{2}=O_{p}\{n^{-bc/(bc+1)}\}$ using regularization $\lambda=n^{-b/(bc+1)}$, when $b\in(1,\infty)$, $c\in(1,2]$, and $\bar{M}$ is fixed. This rate is optimal for some choices of $m$ when Assumption 3(ii) also holds. See Section 5 for details, including results when $b=\infty$, $c=1$, or $\bar{M}_{h}\uparrow\infty$.
Corollary 1 (Main corollary).
Under weak regularity conditions and correct specification of $\hat{\gamma}$ and $\hat{\alpha}$, perhaps outside of the RKHS, if $\bar{M}$ is fixed then $\hat{\theta}$ converges to $\theta_{0}$ at the rate $n^{-1/2}$ and is asymptotically normal. Under mis-specification of $\hat{\gamma}$ or $\hat{\alpha}$, $\hat{\theta}$ converges to $\theta_{0}$, albeit at a slower rate. For pointwise approximations of nonparametric causal functions with $\bar{M}_{h}\uparrow\infty$, $\hat{\theta}$ converges to $\theta_{0}$ at the rate $(nh)^{-1/2}$ and is asymptotically normal when the bandwidth is $h=o(1)$ and $n^{-1/2}h^{-3/2}=o(1)$. See Section 5 and Appendix A.
Proposition 1 (A practical result).
@@ 4.1   General loss function
An important departure from many works on kernel ridge balancing weights is that I allow incorrect specification by the RKHS. [^6] See Remark 3 below. In particular, I allow $\gamma_{0}\not\in H$ and $\alpha_{0}\not\in H$. For the scenario where $\alpha_{0}\not\in H$, I define the best kernel approximation $\alpha_{H}\in H$ to the Riesz representer $\alpha_{0}\in L_{2}$. I defer proofs for this section to Appendix B.
Definition 1 (Kernel approximation to Riesz representer).
@@
Assumption 4 (Original spaces).
@@
Assumption 5 (RKHS regularity).
@@
The space $\mathcal{W}$ accommodates general data types, e.g. texts, images, and graphs. Commonly used kernels are bounded. Boundedness implies that the counterfactual kernel mean embedding $\mu^{(m)}=\mathbb{E}\{\phi^{(m)}(W)\}$ and the uncentered covariance operator $T=\mathbb{E}\{\phi(W)\otimes\phi(W)^{*}\}$ are well defined. [^7] Here, I use outer product notation: $(a\otimes b^{*})c=a\langle b,c\rangle_{H}$, where $b^{*}$ is the adjoint of $b$. This counterfactual kernel mean embedding generalizes the well known kernel mean embedding [Smola et al., 2007]. It is unconditional and does not involve the outcome. Measurability is another standard condition.
The characteristic property ensures that no counterfactual information is lost when approximating $\alpha_{0}$ as $\alpha_{H}$. In particular, it means that $\mathbb{P}\mapsto\mu^{(m)}$ is injective, so $\mu^{(m)}$ may be used as a sufficient statistic for $\alpha_{H}$, as shown in the following lemma.
Lemma 3 (Towards a loss for KRRR).
Suppose Assumptions 14, and 5 hold. Then we have $\alpha_{H}\in\operatorname*{\arg\!\min}_{\alpha\in H}\mathcal{L}(\alpha)$ where $\mathcal{L}(\alpha)=-2\langle\alpha,\mu^{(m)}\rangle_{H}+\langle\alpha,T\alpha\rangle_{H}$.
The sufficient statistics $\mu^{(m)}\in H$ and $T\in H\otimes H^{*}$ are infinite dimensional generalizations of the counterfactual moment and the covariance matrix in automatic debiased machine learning (Auto-DML) [Chernozhukov et al., 2022b, Chernozhukov et al., 2022a]. Auto-DML uses $p$ explicit basis functions, and its sufficient statistic are a vector in $\mathbb{R}^{p}$ and a matrix in $\mathbb{R}^{p\times p}$. Kernel balancing weights use the countable eigenfunctions of the kernel $k$ as infinitely many implicit basis functions.
Inspired by Lemma 3, I study a general loss for kernel ridge Riesz representers (KRRR) at the population level. As we will see below, its empirical analogue recovers the loss for kernel ridge balancing weights previously proposed for treatment effects. Let $\hat{\mu}^{(m)}=\mathbb{E}_{n}\{\phi^{(m)}(W)\}$ and $\hat{T}=\mathbb{E}_{n}\{\phi(W)\otimes\phi(W)^{*}\}$ be the empirical summary statistics.
Definition 2 (Loss for KRRR).
@@
The KRRR loss in Definition 2 differs from Dantzig [Chernozhukov et al., 2022b], Lasso [Chernozhukov et al., 2022a], and adversarial [Chernozhukov et al., 2020] losses for Riesz representers whose generalization errors have been previously analyzed. The KRRR loss has a simple first order condition which implies a nonasymptotic balance-variance trade-off similar to those of kernel balancing weights in the literature [Hazlett, 2020, Wong and Chan, 2018, Zhao, 2019, Kallus, 2020, Hirshberg et al., 2019].
Lemma 4 (First order condition).
Suppose the conditions of Lemma 3 hold. The first order conditions give $T\alpha_{H}=\mu^{(m)}$, $\alpha_{\lambda}=(T+\lambda)^{-1}\mu^{(m)}$, and $\tilde{\alpha}=(\hat{T}+\lambda)^{-1}\hat{\mu}^{(m)}.$
Corollary 2 (Balance-variance trade-off).
Suppose the conditions of Lemma 3 hold. Then we have $\|\mathbb{E}_{n}\{\tilde{\alpha}(W)\phi(W)-\phi^{(m)}(W)\}\|_{H}/\|\tilde{\alpha}\|_{H}=\lambda$.
By Definition 2, a smaller regularization value $\lambda$ translates to an estimator $\tilde{\alpha}$ with a smaller bias and a larger variance. Corollary 2 illustrates that a smaller (normalized) imbalance of the features, $\|\mathbb{E}_{n}[\tilde{\alpha}(W)\phi(W)-\phi^{(m)}(W)]\|_{H}/\|\tilde{\alpha}\|_{H}$, is possible if and only if we pay the cost of a larger variance. In the finite sample analysis to follow, $\lambda\downarrow 0$ in a way that optimally navigates this trade-off (when rate lower bounds are known). In particular, we will not require undersmoothing for inference on the causal parameter.
4.2   Detailed comparisons: Ridge regression and balancing weights
Lemma 4 has several more corollaries, which formalize the sense in which KRRR builds on and extends previous frameworks. In these corollaries, let $\tilde{\gamma}$ be the kernel ridge regression estimator. Throughout, I assume that the conditions of Lemma 3 hold.
Corollary 3 (Special case: Kernel ridge regression).
@@ 4.2   Detailed comparisons: Ridge regression and balancing weights
Definition 3 (Loss for kernel ridge balancing weights).
@@
Corollary 4 (Special case: Kernel ridge balancing weights; eq. 8 of [Bruns-Smith et al., 2023]).
On the training data, $\tilde{\alpha}(W_{i})=\tilde{\beta}_{i}$. Appendix D.3 of [Hirshberg et al., 2019] studies $\frac{1}{n}\sum_{i=1}^{n}\{\tilde{\beta}_{i}-\alpha_{0}(D_{i},X_{i})\}^{2}$, i.e. training error.
In summary, KRRR nests kernel ridge regression and kernel ridge balancing weights, unlike the adversarial Riesz representer of [Chernozhukov et al., 2020]. The connection to kernel ridge regression foreshadows the new generalization error guarantees for KRRR.
Next, I recap known equivalences for the estimation of treatment effects [Speckman, 1979, Kallus, 2020, Hirshberg et al., 2019] and more generally for the estimation of causal parameters within the class of Assumption 1 [Bruns-Smith et al., 2023]. With $\hat{\alpha}$ taken to be KRRR $\tilde{\alpha}$, these equivalences only hold when the regression estimator $\hat{\gamma}$ is kernel ridge regression $\tilde{\gamma}$, and when the same observations are used for $\tilde{\gamma}$, $\tilde{\alpha}$, and the causal parameter.
Corollary 5 (Equivalence in a special case; Proposition 27 of [Kallus, 2020]).
@@
Corollary 6 (Non-equivalence in general).